# TDSM: Triplet Diffusion for Skeleton-Text Matching in Zero-Shot Action Recognition

### v1

@@ 3.4 Inference Phase
Importantly, human actions often share common skeletal movements across related actions, and our model leverages them within seen skeleton data to extrapolate and associate unseen skeleton sequences with their corresponding text prompts. This capability to infer from familiar patterns reinforces the model’s discriminative power, enabling reliable zero-shot recognition of previously unseen actions.
In inference, each unseen skeleton sequence $\mathbf{X}^{u}$ and its all candidate text prompts are inputted to our TDSM, and the resulting noises for the all candidate text prompts are compared with a fixed GT noise $\bm{\epsilon}_{\text{test}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $\mathbf{X}^{u}$ is first encoded into the skeleton latent space through $\mathcal{E}_{x}$ as:
@@
$\hat{\bm{\epsilon}}_{k}=\mathcal{T}_{\text{diff}}(\mathbf{z}_{x,t}^{u},\mathbf{z}_{g,k}^{u},\mathbf{z}_{l,k}^{u},t_{\text{test}}).$ (18)
@@
This process ensures that the model selects the action label whose text prompt well aligns with the skeleton sequence, enabling accurate zero-shot action recognition for unseen skeleton sequences with unseen action labels.
@@ 4.1 Datasets
Table 1: Top-1 accuracy results of various zero-shot skeleton-based action recognition methods evaluated on the SynSE and PURLS benchmarks for the NTU-60 and NTU-120 datasets. Each split is denoted as X/Y, where X represents the number of seen classes and Y the number of unseen classes. The results in red highlight the best-performing model, while those in blue indicate the second-best. For our TDSM framework, the reported accuracy is the average value obtained from 10 trials, each with different Gaussian noise.
@@ 4.2 Experiment Setup
Our TDSM were implemented in PyTorch [40] and conducted on a single NVIDIA GeForce RTX 3090 GPU. Our model variants were trained for 50,000 iterations, with a warm-up period of 100 steps. We employed the AdamW optimizer [38] with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.01. A cosine-annealing scheduler [37] was used to dynamically update the learning rate at each iteration. The batch size was set to 256, but for the TD loss computation, the batch was duplicated for positive and negative samples, resulting in an effective batch size of 512. Through empirical validation, the loss weight $\lambda$ and margin $\tau$ were both set to 1.0. The diffusion process was trained with a total timestep of $T=50$. For inference, the optimal timestep $t_{\text{test}}=25$ was selected based on accuracy trends across datasets (Fig. 5). We utilized $B=12$ CrossDiT Blocks, each containing a multi-head self-attention module with 12 heads. All feature dimensions were set to $C=768$. The local text features contained $M_{l}=35$ tokens, while the skeleton features were represented by a single token $M_{x}=1$. To ensure reproducibility, the random seed was fixed at 2,025 throughout all experiments. For the SynSE [16] and PURLS [71] seen and unseen split settings, we used the Shift-GCN [8] architecture as our skeleton encoder. In the SMIE [69] split setting, we employed the ST-GCN [63] structure. Additionally, for the SynSE and SMIE settings, we utilized both the text prompts and skeleton features provided by SA-DVAE [32]. However, as the PURLS setting did not offer pre-trained skeleton features, we directly trained them using cross-entropy loss $\mathcal{L}_{\text{CE}}$. Across all tasks, we adopted the text encoder from CLIP-ViT-H/14 [44, 21] to transform text prompts into latent representations.
@@ 4.3 Performance Evaluation
Table 2: Top-1 accuracy results of various zero-shot skeleton-based action recognition methods evaluated on the NTU-60, NTU-120, and PKU-MMD datasets under the SMIE benchmark. The reported values are the average performance across three splits.
@@ 4.4 Ablation Studies
Table 3: Top-1 accuracy of different loss function configurations on the NTU-60 and NTU-120 datasets. The results compare models trained with only the diffusion loss $\mathcal{L}_{\text{diff}}$, only the triplet diffusion loss $\mathcal{L}_{\text{TD}}$, and the combination of both losses.
@@
Table 4: Top-1 accuracy with varying total timesteps $T$ on the NTU-60 and NTU-120 datasets.

### v3

@@ 4.4 Inference Phase
In inference (Fig. 3), each unseen skeleton sequence $\mathbf{X}^{u}$ and its all candidate text prompts are inputted to our TDSM, and the resulting noises for the all candidate text prompts are compared with a fixed GT noise $\bm{\epsilon}_{\text{test}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $\mathbf{X}^{u}$ is first encoded into the skeleton latent space through $\mathcal{E}_{x}$ as:
@@
$\hat{\bm{\epsilon}}_{k}=\mathcal{T}_{\text{diff}}(\mathbf{z}_{x,t}^{u},t_{\text{test}};\mathbf{z}_{g,k}^{u},\mathbf{z}_{l,k}^{u}).$ (14)
@@
This process ensures that the model selects the action label whose text prompt well aligns with the skeleton sequence, enabling accurate zero-shot action recognition for unseen skeleton sequences with unseen action labels. Unlike the generative models that iteratively refine samples, our TDSM performs a one-step inference at a fixed timestep, making it efficient and well-suited for discriminative skeleton-text alignment.
Figure 3: Inference framework of our TDSM for ZSAR.
@@ 5.1 Datasets
Table 1: Top-1 accuracy results of various zero-shot skeleton-based action recognition (ZSAR) methods evaluated on the SynSE and PURLS benchmarks for the NTU-60 and NTU-120 datasets. Each split is denoted as X/Y, where X represents the number of seen classes and Y the number of unseen classes. The results in red highlight the best-performing model, while those in blue indicate the second-best. For our TDSM framework, the reported accuracy is the average value obtained from 10 trials, each with different Gaussian noise.
@@ 5.2 Experiment Setup
Our TDSM were implemented in PyTorch [31] and conducted on a single NVIDIA GeForce RTX 3090 GPU. Our model variants were trained for 50,000 iterations, with a warm-up period of 100 steps. We employed the AdamW optimizer [30] with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.01. A cosine-annealing scheduler [29] was used to dynamically update the learning rate at each iteration. The batch size was set to 256, but for the TD loss computation, the batch was duplicated for positive and negative samples, resulting in an effective batch size of 512. Through empirical validation, the loss weight $\lambda$ and margin $\tau$ were both set to 1.0. The diffusion process was trained with a total timestep of $T=50$. For inference, the optimal timestep $t_{\text{test}}=25$ was selected based on accuracy trends across datasets (Fig. 4). For the SynSE [14] and PURLS [53] seen and unseen split settings, we used the Shift-GCN [7] architecture as our skeleton encoder. In the SMIE [51] split setting, we employed the ST-GCN [47] structure. We adopted Shift-GCN and ST-GCN because the majority of previous ZSAR studies use them, ensuring a fair comparison without a confounding factor of potentially more powerful encoders. For fair comparison, we used the same text prompts employed in existing works. Across all tasks, we adopted the text encoder from CLIP [34, 19] to transform text prompts into latent representations.
@@ 5.3 Performance Evaluation
Table 2: Top-1 accuracy results of various ZSAR methods evaluated on the NTU-60, NTU-120, and PKU-MMD datasets under the SMIE benchmark. The reported values are the average performance across three splits.
@@ 5.4 Ablation Studies
Table 3: Ablation study on loss function configurations. The results compare models trained with only the diffusion loss $\mathcal{L}_{\text{diff}}$, only the triplet diffusion loss $\mathcal{L}_{\text{TD}}$, and their combination.
Contribution of global and local text features. Our TDSM framework employs two types of text features for skeleton-text matching: a global text feature $\mathbf{z}_{g}$ that encodes the entire sentence as a single token, and local text features $\mathbf{z}_{l}$ that preserve token-level details for each word in the sentence. As shown in Table 4, combining the global and local text features achieves the best performance, outperforming the models that use either feature independently. The global text feature provides overall discriminative power by capturing high-level semantics of the action description, enabling robust matching across diverse action categories. Meanwhile, the local text features retain finer details that are effective for distinguishing subtle differences between semantically similar actions.
  Global $\mathbf{z}_{g}$ Local $\mathbf{z}_{l}$ NTU-60 (Acc, %) NTU-120  (Acc, %)
55/5 split 48/12 split 110/10 split 96/24 split
83.41 51.50 70.14 61.90
83.33 52.63 69.95 62.10
86.49 56.03 74.15 65.06
Table 4: Ablation study on text feature types. The results compare models trained with only $\mathbf{z}_{g}$, only $\mathbf{z}_{l}$, and their combination.
@@
Effect of random Gaussian noise. To examine the role of Gaussian noise in training, we conducted ablation studies by using fixed Gaussian noise during both training and inference, instead of introducing new random noise at each training step. As shown in Table 6, using fixed Gaussian noise overly simplifies the learning process, causing the network to overfit specific noise patterns and reducing its generalization ability. In contrast, introducing random Gaussian noise at each step increases variability in the learning process, acting as a regularization mechanism that prevents overfitting. This enhances model robustness and improves the alignment between skeleton features and text prompts.
@@
Table 5: Ablation study on the impact of total timesteps $T$ in the training of the diffusion process.
  Gaussian noise $\bm{\epsilon}$ NTU-60 (Acc, %) NTU-120  (Acc, %)
55/5 split 48/12 split 110/10 split 96/24 split
Fixed 76.40 44.25 64.01 52.21
Random 86.49 56.03 74.15 65.06
Table 6: Ablation study on the effect of noise $\bm{\epsilon}$ during training.